OpenStudy (anonymous):

Pleaase help...What is the difference between sin x+4Π and sin(x+4Π )? I need to find which one equals "sin x"

OpenStudy (anonymous):

sin x+4"pie" would be true, correct?

OpenStudy (anonymous):

true for what?

OpenStudy (anonymous):

The question asks which equation is "true"... one of these equals "sin x"

OpenStudy (campbell_st):

think about the graph y = x has a y intercept y 0 y = x + 2 has a y intercept of y = 2 its a parallel line and moved "up" the y axis 2 units y = x - 5 has a y-intercept y = -5... parallel to y = x but moved"down" the y axis 5 units the same principal can be applied to y = sin(x) + 4pi... same graph as y = sin(x) just moved up............ units y = sin(x + 4pi) is saying add 2 revolutions to x then find the sin... or in terms of degrees x + 720.... so do the quadrants for the unit circle change hope it makes sense...

OpenStudy (anonymous):

I think i get this.. so you are saying sin x +4pi does NOT equal sin x?

OpenStudy (campbell_st):

its the graph but translated....

OpenStudy (anonymous):

Right and I need to find the one where placed on top of one another, the graphs are identical.. so that would not be it...